OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 9?

OpenStudy (anonymous):

The standard equation for a circle is \[(x - h)^{2} + (y - k)^{2} = r ^{2}\] with (h,k) as your center and r being the radius.

OpenStudy (anonymous):

a) x^2 + y^2 = 3 b) x^2 + y^2 = 9 c) x^2 + y^2 = 81